OpenStudy (anonymous):

A waitress earns an average tip of 12% of the cost of the food she serves. If she serves $375 worth of food in one evening, how much money in tips will she earn on average?

OpenStudy (anonymous):

$375*12% = 375 * 0.12 = 45

OpenStudy (anonymous):

12%=.12 375*.12=45 $45